OpenStudy (anonymous):

y = (cos 5x)^x find y'

OpenStudy (anonymous):

@campbell_st, the power rule doesn't work for variable powers. Take the logarithm of \(y\) first: \[\begin{align*} y&=(\cos5x)^x\\\\ \ln y&=x\ln(\cos 5x)\\\\ \frac{dy}{dx}[\ln y(x)]&=\frac{d}{dx}[x]\ln(\cos5x)+x\frac{d}{dx}[\ln(\cos5x)]\\\\ \frac{1}{y}\frac{dy}{dx}&=\ln(\cos5x)+x\frac{\dfrac{d}{dx}[\cos5x]}{\cos5x}\\\\ \frac{1}{(\cos5x)^x}\frac{dy}{dx}&=\ln(\cos5x)-5x\frac{\sin5x}{\cos5x}\\\\ y'&=(\cos5x)^x\left(\ln(\cos5x)-5x\tan5x\right) \end{align*}\]